OpenStudy (anonymous):

Calculus help? A company introduces a new product for which the number of units sold, S, is projected to be S(t)= 200(5- (9/2+ t)) where t is the time in months. a)Find the mean value of S(t) during the first year using 0 for the start year and 12 for the end. b) During what month does S'(t) equal the mean value?

OpenStudy (dumbcow):

\[avg = \frac{1}{b-a} \int\limits_a^b S(t) dt\]

OpenStudy (dumbcow):

S'(t) = -200

OpenStudy (anonymous):

is that the mean? I don't understand...

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@phi

OpenStudy (dumbcow):

sorry i was away...yes avg is same as mean are you sure you wrote function correctly? it doesn't make sense for the derivative to be a constant

OpenStudy (anonymous):

well the function is set up \[S(t)= 200 (5 - \frac{9 }{ 2+t })\]

OpenStudy (dumbcow):

ahh ok i misread it

OpenStudy (dumbcow):

so you want to set mean = derivative \[\frac{1}{12} \int\limits_0^{12} S(t) dt = S'(t)\]

OpenStudy (dumbcow):

for integral use fact that \[\int\limits \frac{1}{u} = \ln u\]

OpenStudy (anonymous):

I have no idea how to solve that...........

OpenStudy (dumbcow):

ok \[\frac{50}{3} \int\limits_0^{12} (5 - \frac{9}{2+t}) dt\] \[= \frac{50}{3} |_0^{12} 5t -9 \ln(2+t)\] \[=\frac{50}{3} [ (60-9 \ln(14)) - (0-9\ln(2)) ]\] \[= 1000 - 150(\ln(14) -\ln(2))\] \[=1000 - 150 \ln 7\]

OpenStudy (anonymous):

thats the last answer because there are no like terms right?

OpenStudy (dumbcow):

right so that is mean (part a), now you set that equal to derivative for part b

OpenStudy (anonymous):

the derivative is s'(t)= 1800/(t+2)^2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so 1000-150In7 = 1800/(t+2)^2 ?

OpenStudy (anonymous):

and how do I solve that?

OpenStudy (dumbcow):

use algebra to isolate "t" first i would move (t+2)^2 to other side by multiplying

OpenStudy (dumbcow):

to help simplify you could replace"1000 -150ln7" with a constant like "k" \[k = \frac{1800}{(t+2)^{2}}\]

OpenStudy (dumbcow):

\[(t+2)^{2} = \frac{1800}{k}\]

OpenStudy (dumbcow):

something is wrong the answer is a neg "t" value. are you learning about the Mean Value Thm ?? i think you want the mean value of S'(t) ??

OpenStudy (anonymous):

sorry guys just came back

OpenStudy (anonymous):

they said S(t) not the derivative, I tried wolframalpha but it doesn't undrestand the inquiry

OpenStudy (anonymous):

oh is that how you type it in wolframalpha?

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

its seems pretty big number though?,,

OpenStudy (anonymous):

yeah I see that. ok so how do i find b

OpenStudy (loser66):

@jdoe0001 @jigglypuff314

OpenStudy (anonymous):

I don't get it. the problem is wrong?